# KunLunBaizeRAG: Reinforcement Learning Driven Inference Performance Leap for Large Language Models

### v1

@@ 3.2 Main Results
The experimental results demonstrate that our method significantly outperforms other baseline models across all datasets when compared to the new data. Taking the HotpotQA dataset as an example: KunLunBaizeRAG - Baize - 7B-instruct achieves an EM score of 41.52 and an LJ score of 61.62. This represents a substantial improvement over Naive Generation, which only attains an EM score of 23.63 and an LJ score of 36.26. Compared to Naive RAG, which achieves an EM score of 34.96 and an LJ score of 54.73, KunLunBaizeRAG - Baize - 7B shows an improvement of 6.56 in EM and 6.89 in LJ. Notably, when compared to the recent Search - o1 model, KunLunBaizeRAG - Baize - 7B-instruct also exhibits superior performance. Search - o1 achieves an EM score of 37.64 and an LJ score of 56.15 on the HotpotQA dataset. In contrast, KunLunBaizeRAG - Baize - 7B-instruct improves the EM score by 3.88 and the LJ score by 5.47. The data fully demonstrates the effectiveness of our method in complex reasoning tasks. Through the innovative RAG - Driven Reasoning Alignment (RDRA) and Search - Think Iterative Enhancement (STIE) mechanisms, as well as the newly introduced NLR intelligent routing mechanism, our model achieves remarkable performance improvements. Moreover, our method shows robust performance across different model scales and can dynamically balance local and network knowledge, which further enhances the practicality and broad applicability of this method in various complex reasoning tasks. It provides a new and effective solution for the application of large language models in complex reasoning tasks, holding significant research value and application potential.
@@
KunLunBaizeRAG - Baize - 7B 39.37 58.76 37.67 48.56 20.28 30.59 41.21 51.43
KunLunBaizeRAG - Baize - 7B - Instruct 41.52 61.62 38.59 50.22 20.33 31.43 40.42 51.72
@@
KunLunBaizeRAG - Baize - 32B 41.77 63.27 39.52 50.59 25.44 36.57 53.44 65.46
KunLunBaizeRAG - Baize - 32B - Instruct 43.73 64.75 40.94 51.39 26.14 37.56 54.86 66.28
@@ 3.4 Case Study
<|im_start|> system You are a helpful assistant that can solve the given question step by step with the help of the wikipedia search tool. Given a question, you need to first think about the reasoning process in the mind and then provide the answer. During the thinking process, if necessary, you can call the Wikipedia search tool to search for factual information on specific topics. Of course, the user may provide some relevant background information. If the answer to the question is clearly contained in the relevant background provided by the user, you can directly provide the answer without calling the Wikipedia search tool. You need to determine the length of the final answer according to the user’s question. For simple questions, the more concise the answer is, the better, and for complex questions, the more detailed the answer is, the better. The reasoning process and answer are enclosed within ¡think¿ ¡/think¿ and ¡answer¿ ¡/answer¿ tags respectively, and the search query and result are enclosed within ¡search¿ ¡/search¿ and ¡result¿ ¡/result¿ tags respectively. For ¡answer¿, our model uses different answer templates for different tasks. It uses ¡short_answer¿ to answer specific questions and ¡long_answer¿ to answer summary questions. For example:
@@
In the last part of the answer, the final exact answer is enclosed within boxed with latex format.¡—im_end—¿¡—im_start—¿user
@@
To gain a more intuitive understanding of KunLunBaizeRAG’s utility, we present a case study in Table 3 from the Baize-32B-Instruct model’s reinforcement learning process. When faced with the question ”Which film came out first, Blind Shaft or The Mask Of Fu Manchu?” the model first analyzes the content in the ¡background¿ tag to extract and organize background information related to the two films. Although the ¡background¿ tag does not directly mention their release dates, it provides some context about the films’ genres and styles, helping the model better position its search.
Next, the model delves into the problem-solving approach in the ¡think¿ tag and realizes that the key to determining the release order of the films is to obtain their exact release years. Therefore, the model constructs precise search requests within the ¡search¿ tag, such as ”Blind Shaft release date” and ”The Mask Of Fu Manchu release date,” to obtain clear release date information from the retrieval environment.
After the retrieval results (the content of the ¡result¿ tag) are returned, the model combines the background knowledge previously gained from the ¡background¿ tag to conduct a comprehensive analysis of the search results. For instance, if the ¡background¿ tag mentions some historical context or genre characteristics of the films, the model can use this information to assess the credibility of the search results or determine if further verification is needed.
Finally, within the ¡answer¿ tag, the model provides an accurate response based on the detailed reasoning process and verification results. Throughout this case, it is evident that the model not only relies on search tools to gather information but also fully utilizes the background knowledge from the ¡background¿ tag to aid understanding and verification. Additionally, the clarity of the question (the ¡question¿ tag) guides the model’s thinking and search direction. This demonstrates the model’s ability to perform in-depth reasoning and provide accurate answers to complex questions under the collaboration of multiple tags.

### v2

@@ 3.2 Main Results
The experimental results demonstrate that our method significantly outperforms other baseline models across all datasets when compared to the new data. Taking the HotpotQA dataset as an example: KunLunBaizeRAG - Baize - 7B-instruct(KLBRAG-Baize-7B-Ins) achieves an EM score of 41.52 and an LJ score of 61.62. This represents a substantial improvement over Naive Generation, which only attains an EM score of 23.63 and an LJ score of 36.26. Compared to Naive RAG, which achieves an EM score of 34.96 and an LJ score of 54.73, KunLunBaizeRAG - Baize - 7B shows an improvement of 6.56 in EM and 6.89 in LJ. Notably, when compared to the recent Search - o1 model, KunLunBaizeRAG - Baize - 7B-instruct also exhibits superior performance. Search - o1 achieves an EM score of 37.64 and an LJ score of 56.15 on the HotpotQA dataset. In contrast, KunLunBaizeRAG - Baize - 7B-instruct improves the EM score by 3.88 and the LJ score by 5.47. The data fully demonstrates the effectiveness of our method in complex reasoning tasks. Through the innovative RAG - Driven Reasoning Alignment (RDRA) and Search - Think Iterative Enhancement (STIE) mechanisms, as well as the newly introduced NLR intelligent routing mechanism, our model achieves remarkable performance improvements. Moreover, our method shows robust performance across different model scales and can dynamically balance local and network knowledge, which further enhances the practicality and broad applicability of this method in various complex reasoning tasks. It provides a new and effective solution for the application of large language models in complex reasoning tasks, holding significant research value and application potential.
@@
KLBRAG-Baize-7B 39.37 58.76 37.67 48.56 20.28 30.59 41.21 51.43
KLBRAG-Baize-7B-Ins 41.52 61.62 38.59 50.22 20.33 31.43 40.42 51.72
@@
KLBRAG-Baize-32B 41.77 63.27 39.52 50.59 25.44 36.57 53.44 65.46
KLBRAG-Baize-32B-Ins 43.73 64.75 40.94 51.39 26.14 37.56 54.86 66.28
@@ 3.4 Case Study
<|im_start|> system You are a helpful assistant that can solve the given question step by step with the help of the wikipedia search tool. Given a question, you need to first think about the reasoning process in the mind and then provide the answer. During the thinking process, if necessary, you can call the Wikipedia search tool to search for factual information on specific topics. Of course, the user may provide some relevant background information. If the answer to the question is clearly contained in the relevant background provided by the user, you can directly provide the answer without calling the Wikipedia search tool. You need to determine the length of the final answer according to the user’s question. For simple questions, the more concise the answer is, the better, and for complex questions, the more detailed the answer is, the better. The reasoning process and answer are enclosed within <$\textless think\textgreater$ > <$>\textless/think\textgreater$ and <$\textless answer\textgreater$ > <$>\textless/answer\textgreater$ tags respectively, and the search query and result are enclosed within <$\textless search\textgreater$ > <$>\textless/search\textgreater$ and <$\textless result\textgreater$ > <$>\textless/result\textgreater$ tags respectively. For <$\textless answer\textgreater$ >, our model uses different answer templates for different tasks. It uses <$\textless short\_answer\textgreater$ > to answer specific questions and <$\textless long\_answer\textgreater$ > to answer summary questions. For example:
@@
In the last part of the answer, the final exact answer is enclosed within boxed with latex format.<$\textless|im\_end\textgreater$ > <$\textless|im\_start\textgreater$ >user
@@
To gain a more intuitive understanding of KunLunBaizeRAG’s utility, we present a case study in Table 3 from the Baize-32B-Instruct model’s reinforcement learning process. When faced with the question ”Which film came out first, Blind Shaft or The Mask Of Fu Manchu?” the model first analyzes the content in the <$\textless background\textgreater$ > tag to extract and organize background information related to the two films. Although the ¡background¿ tag does not directly mention their release dates, it provides some context about the films’ genres and styles, helping the model better position its search.
Next, the model delves into the problem-solving approach in the <$\textless think\textgreater$ > tag and realizes that the key to determining the release order of the films is to obtain their exact release years. Therefore, the model constructs precise search requests within the <$\textless search\textgreater$ > tag, such as ”Blind Shaft release date” and ”The Mask Of Fu Manchu release date,” to obtain clear release date information from the retrieval environment.
After the retrieval results (the content of the <$\textless result\textgreater$ > tag) are returned, the model combines the background knowledge previously gained from the <$\textless background\textgreater$ > tag to conduct a comprehensive analysis of the search results. For instance, if the ¡background¿ tag mentions some historical context or genre characteristics of the films, the model can use this information to assess the credibility of the search results or determine if further verification is needed.
Finally, within the <$\textless answer\textgreater$ > tag, the model provides an accurate response based on the detailed reasoning process and verification results. Throughout this case, it is evident that the model not only relies on search tools to gather information but also fully utilizes the background knowledge from the <$\textless background\textgreater$ > tag to aid understanding and verification. Additionally, the clarity of the question (the <$\textless question\textgreater$ > tag) guides the model’s thinking and search direction. This demonstrates the model’s ability to perform in-depth reasoning and provide accurate answers to complex questions under the collaboration of multiple tags.